OpenStudy (anonymous):

please help find the variable 5x+12=3x+128

OpenStudy (anonymous):

baby im going to help u

OpenStudy (johnweldon1993):

So you ultimately want to solve for 'x' \[\large 5x + 12 = 3x + 128\] Combine like terms "that is to say...combine the variables...subtract 3x from both sides of this equation" 5x + 12 = 3x + 128 -3x -3x -------------------- 2x + 12 = 128 Now you want to get that 'x' alone...so subtract 12 from both sides 2x + 12 = 128 - 12 -12 -------------- 2x = 116 And finally to solve for 'x' you divide both sides by 2 2x = 116 --------- 2 2 \[\large x = \frac{116}{2} = \space ?\]

OpenStudy (anonymous):

there u lovestar